# Coupled micro-resonators for second-order integrated nonlinear optics

### v1

@@ Abstract
for $p\in[0,3]_{\mathbb{N}}$ with $\Delta\beta=\beta_{2}-2\beta_{1}$; $s_{0}=0$ for $p=\{0,2\}$ and $s_{0}=L$ for $p=\{1,3\}$ Defining $\varphi_{2}=-\beta_{2}L$, we obtain the expression of the amplitude of the output field on the trough port
@@
We apply the previous model to the case of two identical coupled GaP microdisks grown along the $z=[001]$ crystallographic axis and embedded in air as described in Fig. 2. The disks are separated by an air gap $g$, their thickness is $H=136~{}\mathrm{nm}$ and their radius $R=2.7~{}\mu\mathrm{m}$.
@@
Thus, the detuning $\delta$ can be calculated varying the air-gap as shown in Fig. 4.a). The corresponding coupling coefficients necessary to calculate the SHG in the coupled mode framework using Eq. (8) are deduced from $|\kappa_{1}|=\sin{\left(\frac{\pi\Delta\nu_{F}}{\Delta_{1}}\right)}$ and $|\kappa_{2}|=\sin{\left(\frac{\pi\Delta\nu_{SH}}{\Delta_{2}}\right)}$ where $\Delta_{1}$ an $\Delta_{2}$ are respectively the free spectral range (FSR) for the F and SH fields [31], the resulting values are plotted in Fig. 4.b). For a microdisk, the effective nonlinear susceptibilities is calculated for the single resonator WGM modes of resonant frequencies $\nu_{F0}$ and $\nu_{SH0}$ and azimutal quantum numbers $m_{1}=17$ and $m_{2}=36$. Cylindrical coordinates $(r,\theta)$ are defined in Fig. 2 and $d_{14}=28.5~{}\mathrm{pm/V}$ is the nonlinear coefficient of bulk GaP [32]. From the coupling mode theory [11] we find
[…]
@@
Figure 5 represents the normalized conversion efficiency $\eta=\frac{|B_{T}|^{2}}{|A_{in}|^{2}P_{in}}$ for the studied structure in the case $g=g_{0}$. For the doubly resonant frequency we obtain a maximal conversion efficiency of $360\%/\mathrm{W}$ which is consistent with results calculated for small thin microdisks (with same Q-factors) under perfect phase matching [32]. This result shows that even though we obtained the double resonance with two supermodes with different symmetry it does not affect the efficiency of the device. The interest of the proposed method is that the resonator coupling gives an additional free parameter which would allows to use thicker and largest disks which would be less sensitive to fabrication imperfections. If we suppose a tolerance of $\nu_{F0}/Q_{F}$ (with $Q_{F}$ the loaded Q-factor for the F) for the double resonance in frequency it translates in a 10 nm fabrication tolerance for the air-gap. Note that the relative phase between the two coupling coefficients $\kappa_{1}$ and $\kappa_{2}$ has a negligible effect on the conversion efficiency maximum.

### v2

@@ Abstract
for $p\in[0,3]_{\mathbb{N}}$ with $\Delta\beta=\beta_{2}-2\beta_{1}$; $s_{0}=0$ for $p=\{0,2\}$ and $s_{0}=L$ for $p=\{1,3\}$ . Defining $\varphi_{2}=-\beta_{2}L$, we obtain the expression of the amplitude of the output field on the trough port
@@
In standard nonlinear photonic devices, the effective index mismatch between the F and SH modes depends only on the material and modal dispersions. It is independent of the cavity length (or micro-resonator radius). On the contrary, the phase mismatch compensation offered by the $\overline{4}$ $\chi^{(2)}$ tensor symmetry, see (Eq. 12), as well as the maximal splitting of coupled resonators both scale with the free spectral range (FSR). Their cumulated effect thus requires small geometries to effectively cancel the effective index mismatch. We apply the previous model to the case of two identical coupled GaP microdisks grown along the $z=[001]$ crystallographic axis and embedded in air as described in Fig. 2. The disks are separated by an air gap $g$, their thickness is $H=136~\mathrm{nm}$ and their radius $R=2.7~\mu\mathrm{m}$.
@@
Thus, the detuning $\delta$ can be calculated varying the air-gap as shown in Fig. 4.a). Note that to produce the results given Fig. 4, we interpolated the numerical discrete calculations of $\Delta\nu_{F}(g)$ and $\Delta\nu_{SH}(g)$ using decaying exponential functions. The corresponding coupling coefficients necessary to calculate the SHG in the coupled mode framework using Eq. (8) are deduced from $|\kappa_{1}|=\sin{\left(\frac{\pi\Delta\nu_{F}}{\Delta_{1}}\right)}$ and $|\kappa_{2}|=\sin{\left(\frac{\pi\Delta\nu_{SH}}{\Delta_{2}}\right)}$ where $\Delta_{1}$ an $\Delta_{2}$ are respectively the FSR for the F and SH fields [31], the resulting values are plotted in Fig. 4.b). For a microdisk, the effective nonlinear susceptibilities is calculated for the single resonator WGM modes of resonant frequencies $\nu_{F0}$ and $\nu_{SH0}$ and azimutal quantum numbers $m_{1}=17$ and $m_{2}=36$. Cylindrical coordinates $(r,\theta)$ are defined in Fig. 2 and $d_{14}=28.5~\mathrm{pm/V}$ is the nonlinear coefficient of bulk GaP [32]. From the coupling mode theory [11] we find
$\chi_{0\pm}^{(2)}=-\frac{j\gamma_{2}}{4R^{2}N^{2}_{2}}\int_{0}^{R}r^{2}d_{14}f_{\pm}(r)(\mathcal{H}_{r}^{SH})^{*}dr,$ (14)
@@
Figure 5 represents the normalized conversion efficiency $\eta=\frac{|B_{T}|^{2}}{|A_{in}|^{2}\mathcal{P}_{in}}$ (where $\mathcal{P}_{in}$ is the input F power) for the studied structure in the case $g=g_{0}$. For the doubly resonant frequency we obtain a maximal conversion efficiency of $360\%/\mathrm{W}$ which is consistent with results calculated for small thin microdisks (with same Q-factors) under perfect phase matching [32]. This result shows that even though we obtained the double resonance with two supermodes with different symmetry it does not affect the efficiency of the device. The interest of the proposed method is that the resonator coupling gives an additional free parameter which would allows to use thicker and largest disks which would be less sensitive to fabrication imperfections. As with single doubly resonant resonators, the disk radius must be controlled with an accuracy of the order of $R/Q_{F}$. In the example presented, this gives $0.5~\mathrm{nm}$. For practical implementation, it will be necessary to control the resonances independently, using, for example, thermal effects [33]. Another solution could be to use a single corrugated disk and take advantage of the coupling of co- and counter-propagating modes [34]. This would mean that only one resonator would need to be controlled. If we suppose a tolerance of $\nu_{F0}/Q_{F}$ (with $Q_{F}$ the loaded Q-factor for the F) for the double resonance in frequency it translates in a 10 nm fabrication tolerance for the air-gap. Note that the relative phase between the two coupling coefficients $\kappa_{1}$ and $\kappa_{2}$ has a negligible effect on the conversion efficiency maximum.